OpenStudy (anonymous):

Evaluate 9 − 3 • 2 + 4

OpenStudy (anonymous):

7

OpenStudy (anonymous):

9 - 6 + 4 = 9 - 10 = -1

OpenStudy (anonymous):

7

OpenStudy (anonymous):

oh yes, 7.

OpenStudy (anonymous):

It has to be: (9 - 6) + 4 = 3 + 4 = 7.